If a trapezoid has the sides 8,8,8, and 10, how do you find the area?
Answer by Boreal(15235)

The base is 10, and the other 3 sides are 8.
This makes the trapezoid an isosceles one.
The top side is 8, and the bottom side is 1+8+1, where the 1 makes a right triangle at each end whose hypotenuse, 8 is the side of the trapezoid, and the missing side is the altitude.
The Pythagorean Theorem allows you to get the altitude, which is 64=1^2+x^2
x+sqrt (63)
The area of the trapezoid is the center part, 8* the altitude, which is sqrt(63), and two triangles, each of which has base 1 and altitude sqrt (63). You can use the formula A=(1/2)bh for each triangle, but it is just A=bh, because you have 2 triangles. That is the area.
A=(1/2) h (b1+b2) is the other formula you may use.
You have both bases, 10 and 8, and the altitude, sqrt (63)